FOURIER TRANSFORM MICROWAVE AND INFRARED SPECTROSCOPIC INVESTIGATION OF PRO-PIOLACTONE

\underline{ZIQIU CHEN} and JENNIFER VAN WIJNGAARDEN, Department of Chemistry, University of Manitoba, Winnipeg MB R3T 2N2 Canada.

The pure rotational spectrum of the four-membered ester ring propiolactone (C_{3}H_{4}O_{2}) has been measured in a supersonic jet between 7 and 22 GHz using Fourier transform microwave (FTMW) spectroscopy. For the normal isotopologue, a total of 19 a- and b-type transitions have been recorded. Fifteen transitions due to three different ^{13}C isotopologues have also been observed. The microwave spectrum was analyzed to obtain an improved set of ground state rotational constants in comparison to earlier microwave experiments^{a}. The new set of rotational parameters was used to predict the rovibrational band structure of the lowest frequency modes of propiolactone. A total of 12 vibrational band origins have been observed between 400 and 1500 cm^{−1} using the far infrared beamline of the Canadian Light Source coupled to a Bruker IFS125HR spectrometer. The spectra were recorded with a resolution of 0.000969 cm^{−1} and although the intensities of the bands vary, 9 bands are of sufficient quality for complete rovibrational assignment. The progress of the assignment of this rich spectrum will be discussed.

aD. W. Boone, C O. Britt and J. E. Boggs J. Chem. Phys. \underline{43} (1190), 1965.